Over what grade should we find the 10 best students out of 37 in a test with a average of 82 and a standard deviation of 5?
Answer by ewatrrr(24785)

 
Hi,
m= 82, sd = 10, n = 37 (top 10) 0r top 27 percentile
Normsinv(.135) = z = 1.102 = (x - 81)/7
.6128 = (x - 82)/10
6.128 + 82 = 88.128 Over an 89 test score